# FedEx: Expediting Federated Learning over Heterogeneous Mobile Devices by Overlapping and Participant Selection

### v1

@@ Abstract.
ccs: Computing methodologies Distributed artificial intelligence
1. Introduction
@@
Nowadays, the landscape of federated learning (FL) has undergone a significant shift, which has transitioned from its traditional realm within data centers to embrace the capabilities of mobile devices (19). This transition is supported by the continuous evolution of hardware, propelling mobile devices such as the NVIDIA Xavier, Galaxy Note20, iPad Pro, MacBook laptops, and UAVs into possessing increasingly potent on-device computing capabilities tailored for localized model training (12). Adhering to the core FL principle of conducting localized training of deep neural networks (DNNs) while safeguarding the privacy of raw data, the application of FL to mobile devices has ushered in a host of promising possibilities spanning diverse domains including computer vision (22), natural language processing (13) and many other applications such as e-Healthcare (4), voice assistants in smart home (24), and autonomous driving (30). However, in contrast to GPU clusters with Ethernet connections boasting speeds of 50-100 Gbps (29), wireless transmission bandwidth is rather limited(6), resulting in potentially significant delays in gradient transmission. Besides, the varied computational capabilities and channel conditions among mobile devices often result in “straggler” issues, while the non-independent and identically distributed (non-i.i.d.) local data increases the complexity of training to converge. All these factors result in the high latency of FL training over mobile devices, imposing huge burdens on mobile devices and compromising the performance of corresponding applications.
@@
Different from tons of communication/latency efficient FL papers, the concept of overlapping communication and computing, exemplified by delayed gradient averaging (DGA) in (31), tears down the traditional interleaved workflow of local computing and gradient uploading in FL training and magically “disappears” the communication time/phase. The innovative aspect lies in that it enables mobile devices to continue training their local models while transmitting their model updates to FL server, i.e., burying/overlapping the communication phase into continuous local computing one. Since both computing and communication contribute to FL convergence, such a “continuous computing while transmitting” overlapping method helps to reduce the total latency of FL training and may fundamentally eliminate FL’s communication efficiency issue.
@@
Although overlapping based FL has the potential to accelerate FL training, it introduces model staleness problem and may not work well or work at all in heterogeneous environments (i.e., data/computing/wireless heterogeneity). Briefly, since overlapping lets FL clients use the old global model for ongoing local training, the trained local model parameters become stale. To mitigate the model staleness issue, the update correction mechanism can be integrated in overlapping based FL protocols by using stale averaged gradients for homogeneous FL client devices at the huge cost of their memory storage (31). Heterogeneity among clients’ mobile devices makes the memory cost issue more severe and almost diminishes the benefits of overlapping computing and communication. For example, following typical overlapping based FL protocols, fast devices (computing + communications) have to wait for the slowest device (i.e., straggler) to finish an FL training round, and those fast devices, in particular fast computing devices, will undergo many rounds of local training in advance before receiving the new/next round global model for update correction. That will accumulate rounds of gradients to correct, causing significant model staleness, and impose a huge burden on, or even overflow fast mobile devices’ memory. The advanced local training of those fast devices can also drive local models towards their respective local optima, potentially resulting in divergence from the objective of the global model (a.k.a model drift (16)) and hindering FL convergence.
@@
Figure 1. Overlapping deficiencies due to heterogeneity.
To alleviate the straggler issue in heterogeneous environments, researchers have explored to select a subset of devices to participate in FL training over rounds, i.e., FL participation selection (PS) (28; 25; 11; 21). For example, Oort (18), a pioneering FL PS design, jointly considers the statistical utility (i.e., FL learning convergence/accuracy) and global system utility (i.e., training latency per round). Oort selects participating mobile devices that have good contributions to FL learning convergence and are capable of completing local training before a setup latency threshold. However, existing FL PS methods still follow traditional computing and communication interleaving FL training protocol and don’t consider the training speedup brought by continuous computing or overlapping.
@@
  • We develop a FedEx prototype and evaluate its performance with extensive experiments. The experimental results demonstrate that FedEx can remarkably reduce the latency for FL training over heterogeneous mobile devices at limited costs.
2. Background and Motivations
2.1. FL on Mobile Devices
@@
(1) $\min_{w\in\mathbb{R}^{d}}F\left(w\right):=\sum_{n=1}^{N}\pi_{n}F_{n}\left(w\right)\;,$
@@
Table 1. Comp. vs Comm. (Squeezenet@CIFAR10).
Mobile Device Configuration [^2] The transmission data rates displayed in Table 1 represent the averaged achievable data rates for uploading transmissions. Figure 2 employs the same communication settings. $T_{n}^{cp}$ $T_{n}^{cm}$ $T_{n}^{cm}/T_{n}^{cp}$
NVIDIA Xavier w/ Wi-Fi (6.9Mbps) 1.13s 5.54s 4.90
NVIDIA TX2 w/ Wi-Fi (6.0Mbps) 1.35s 6.40s 4.74
Xiaomi 12S w/ LTE (5.0Mbps) 0.84s 7.66s 9.12
2.2. Deficiencies of Overlapping Based FL in Heterogeneous Environments
@@
Figure 2. Performance comparison: homogeneity vs heterogeneity (CNN@MNIST, non-i.i.d. data).
Specifically, following DGA procedure [^1] Please refer to “Algorithm 1 Delayed Gradient Averaging (DGA)” in (31)., the device transmits its local updates when it completes every fixed number of local training iterations, say $K$ iterations. Nevertheless, it may lead to “collision of local update communications” issues, if certain device’s transmission latency is much larger than its computing time consumption as illustrated in Fig. 1(a). Even if the mobile device can be equipped with multiple antennas allowing for simultaneous transmissions of local updates over different rounds, it will transmit the stale local model gradients, i.e., locally trained based on the ($r-1$)-th round global model, for the ($r+1$)-th round updates before it receives the $r$-th round global model for staleness correction. We conducted empirical studies with popular mobile devices and accessing technologies, and the results in Table 1 show that the fast computing and slow transmission situations largely exist in practice, which can trigger the collision of communications for local model updates.
Assuming the overlapping training can proceed without “collision of local model updates” in heterogeneous environments, the memory cost issue for model update correction will be aggravated due to computing and wireless heterogeneity among mobile devices and overlapping model staleness may incur model drift issue due to data heterogeneity, which can degrade or even fail FL convergence. Overlapping based FL requires each device to locally store multiple copies of its old local model updates for update correction, and the memory consumption of the mobile device increases linearly with the staleness level. Different from homogeneous environments, computing and wireless heterogeneity enlarges the time consumption gap between fast devices and the slowest device (i.e., straggler) to complete their local computing and model updates. The enlarged gap directly increases the staleness level of fast devices. When the overlap staleness $S_{n}$ is large, the heavy gradient storage requirement will impose a huge burden on, or even overflow fast mobile devices’ limited memory. Moreover, due to data heterogeneity, the local trained models point to the direction towards the local optima, which may not be consistent with the objective of the aggregated global model, called model drift (16). Thus, a high overlapping staleness level, i.e., performing too much continuous local computing, is detrimental to the learning accuracy and may significantly slow down or fail FL convergence under data heterogeneous scenarios. Fig. 2 shows our empirical studies to compare the testing accuracy of DGA and FedAvg in homogeneous and heterogeneous environments, respectively, which supports our model drift and staleness analysis above.
@@
2.3. Limitations of SOTA PS Methods
To mitigate the straggler issue in heterogeneous environments, various FL PS designs have been proposed to select a subset of devices to participate in FL training over rounds (28; 25; 11; 21). As an advanced FL PS design, Oort (18) novelly considers both the statistical utility (i.e., FL learning convergence/accuracy) and global system utility (i.e., training latency per round). The utility function of device $n$ in Oort can be expressed as follows.
(2) $\small\begin{split}Util(n)=|B_{n}|\sqrt{\frac{1}{|B_{n}|}\sum_{\begin{subarray}{c}k\in{B}_{n}\end{subarray}}Loss(k)^{2}}\times(\frac{T}{t(n)})^{\mathbb{I}(T<t(n))\times\alpha}\;,\end{split}$
where $B_{n}$ represents the local training samples of mobile device $n$, $Loss(k)$ corresponds to the training loss of the $k$-th sample, $t(n)$ signifies the system latency of mobile device $n$, $T$ is the developer-preferred duration for each round, and $\alpha$ serves as the penalty factor. Additionally, $\mathbb{I}(x)$ denotes an indicator function, where $\mathbb{I}(x)=1$ when $x$ is true, and 00 otherwise. Under Eqn. (2), top $K$ devices with the highest utility will be selected for FL training in synchronous manner. Instead of picking devices with top-$K$ utility deterministically over rounds, Oort employs a “temporal uncertainty” mechanism (18) to gradually increase the utility of a device if it has been overlooked for a long time. Thus, low-utility devices still have chances to be selected to avoid biased FL training.
@@ 2.4. Challenges in Overlapping + PS
Secondly, PS utility function doesn’t take the latency reduction brought by overlapping into account. Simply, the overlapping and continuous training changes the way of traditional computing and communication interleaved FL training, which consequently affects the definition of computing latency in a training round and corresponding per round training latency for a candidate mobile device. While the global system utility (18) in Oort’s utility function is based on conventional per round FL training latency, directly using Oort to select the participating mobile devices for overlapping based FL is not appropriate, since it is not aware of overlapping benefits to accelerate FL training convergence.
@@
3. FedEx Design
3.1. FedEx Overview
@@
Figure 3. The sketch of FedEx procedure and testbed.
3.2. Model Updates with Staleness Ceiling
To preclude the collision of communications, alleviate model drift, constrain the memory cost for continuous computing (Sec. 2.2) and synchronize devices’ local model updates in every round (Sec. 2.4), we propose a novel local model updates protocol with staleness ceiling in FedEx. As discussed in Sec. 2, we found that in heterogeneous environments, existing overlapping principle (e.g. DGA) of uploading local gradients after every $K$ computing iterations is the root for the collision of communications and asynchronous model updates, and its endless/non-constrained continuous local computing aggravates model staleness, which results in model drift as well as huge memory consumption for updates correction. So, to address those issues and fit FL training over heterogeneous mobile devices, FedEx redefines the rule of overlapping local model updates and casts a ceiling on staleness/the iteration number of continuous local computing.
@@ 3.2. Model Updates with Staleness Ceiling
3.3. Overlapping Aware PS Utility Function
@@
To address straggler issues for latency reduction per round (Sec. 2.2) and make PS aware of overlapping benefits (Sec. 2.3 and Sec. 2.4), we develop an overlapping aware utility function for FL PS in heterogeneous environments. Inspired by Oort’s utility function (18) in Eqn. (2), the proposed PS utility function aims to jointly improve statistical utility and delay efficiency of FL training, while considering the latency reduction brought by overlapping computing and communication. Thus, we revise global system utility in Eqn. (2) into overlapping latency utility in FedEx, and present the overlapping aware PS utility function as follows.
@@
(3) $\small\begin{split}U(n,r)\!=\!\underbrace{|B_{n}^{r}|\sqrt{\frac{1}{|B_{n}^{r}|}\sum_{\begin{subarray}{c}k\in{B}_{n}^{r}\end{subarray}}\!Loss(k)^{2}}}_{\textbf{Statistical utility}}\!\times\!\underbrace{\left(\frac{1}{(K\!-\!S_{n}^{r-1})\!\cdot\!t_{n}^{r}\!+\!\frac{D_{n}}{R_{n}^{r}}}\right)^{\alpha}}_{\textbf{Overlapping latency utility}},\end{split}\vspace{-3mm}$
@@
where $|B_{n}^{r}|$ is the number of training data samples of mobile device $n$ in the $r$-th round, $t_{n}^{r}$ is the averaged time consumption for one iteration of local computing at mobile device $n$ (7), $D_{n}$ is the model size of client $n$ and $R_{n}^{r}$ is the transmission rate of mobile device $n$ in the $r$-th round. $\frac{D_{n}}{R_{n}^{r}}$ represents the mobile device $n$’s estimated transmission time. Here, $\alpha$ is the scaling coefficient to trade off the statistical utility and overlapping latency utility.
@@
The utility function above consists of two parts: statistical utility and overlapping latency utility. The statistical utility accounts for the estimated learning performance contributions to FL convergence, which is the same as that in Oort (18). The overlapping latency utility has two components: classical computing latency and the model update transmission latency. For classical computing latency in the $r$-th round, mobile device $n$ needs to complete $K$ local training iterations. Regarding the latency reduction by overlapping, device $n$ has already conducted continuous computing for $S_{n}^{r-1}$ iterations in the ($r-1$)-th round. Thus, in the $r$-th round, device $n$ only needs to perform ($K-S_{n}^{r-1}$) iterations for classical computing phase before transmitting its local gradients. As illustrated in (7), given a learning task and the computing capability of mobile device $n$, $t_{n}^{r}$ is relatively fixed in practice and can be estimated with good precision by averaging the latency for the previous iterations. Therefore, device $n$’s classical computing latency in the $r$-th round can be estimated as $(K-S_{n}^{r-1})\cdot t_{n}^{r}$. For transmission latency, given the fact that mobile devices’ wireless communication conditions don’t significantly change within a short period of time, we assume $R_{n}^{r}$ is fixed in the $r$-th round and can be approximated by using $n$’s transmission rate at the beginning of this round.
@@
The proposed overlapping latency utility is aware of overlapping benefits by considering latency reduction $(K-S_{n}^{r-1})\cdot t_{n}^{r}$. Besides, as discussed in Sec. 3.2, the per-round latency in FedEx is determined by the straggler’s training latency, i.e., the slowest device’s classical on-device training time plus its communication time for transmitting local updates. The overlapping latency utility helps to address straggler issues since it will select fast devices with small training latency in the round, i.e., $[(K-S_{n}^{r-1})\cdot t_{n}^{r}+\frac{D_{n}}{R_{n}^{r}}]$. Together with statistical utility, the proposed overlapping aware utility function [^3] Regarding the privacy preservation of statistical utility and overlap latency utility, we leverage the same privacy analysis as presented in Oort (18). will select the mobile devices with good learning contributions and high overlapping latency utility to participate, which can accelerate the FL training over heterogeneous mobile devices [^4] For device $n$ not selected to participate in the $r$-th round, it refrains from any computing and transmission activities. Instead, it inherits the values of $K$ and $S_{n}^{r-1}$ from the $(r-1)$-th round, updates them into $K$ and $S_{n}^{r}$, and wait for the next time to participate. For devices overlooked for a long time, we follow the “temporal uncertainty” mechanism to gradually increase their utilities, in order to avoid biased training in heterogeneous environments. Due to page limits, the duplicated discussions have been omitted here..
3.4. FedEx Overlapping Trigger

### v3

@@ Abstract
I Introduction
@@
Nowadays, the landscape of federated learning (FL) has undergone a significant shift, which has transitioned from its traditional realm within data centers to embrace the capabilities of mobile devices [1]. This transition is supported by the continuous evolution of hardware, propelling mobile devices such as the NVIDIA Xavier, Galaxy Note20, iPad Pro, MacBook laptops, and UAVs into possessing increasingly potent on-device computing capabilities tailored for localized model training [2]. Adhering to the core FL principle of conducting localized training of deep neural networks (DNNs) while safeguarding the privacy of raw data, the application of FL to mobile devices has ushered in a host of promising possibilities spanning diverse domains including computer vision [3], natural language processing [4] and many other applications such as e-Healthcare [5], voice assistants in smart home [6], and autonomous driving [7]. However, in contrast to GPU clusters with Ethernet connections boasting speeds of 50-100 Gbps [8], wireless transmission bandwidth is rather limited, resulting in potentially significant delays in gradient transmission. Besides, the varied computational capabilities and channel conditions among mobile devices often result in “straggler” issues, while the non-independent and identically distributed (non-i.i.d.) local data increases the complexity of training to converge. All these factors result in the high latency of FL training over mobile devices, imposing huge burdens on mobile devices and compromising the performance of corresponding applications.
@@
Different from tons of communication/latency efficient FL papers, the concept of overlapping communication and computing, exemplified by delayed gradient averaging (DGA) in [9], tears down the traditional interleaved workflow of local computing and gradient uploading in FL training and magically “disappears” the communication time/phase. The innovative aspect lies in that it enables mobile devices to continue training their local models while transmitting their model updates to FL server, i.e., burying/overlapping the communication phase into continuous local computing one. Since both computing and communication contribute to FL convergence, such a “continuous computing while transmitting” overlapping method helps to reduce the total latency of FL training and may fundamentally eliminate FL’s communication efficiency issue.
@@
Although overlapping based FL has the potential to accelerate FL training, it introduces model staleness problem and may not work well or work at all in heterogeneous environments (i.e., data/computing/wireless heterogeneity). Briefly, since overlapping lets FL clients use the old global model for ongoing local training, the trained local model parameters become stale. To mitigate the model staleness issue, the update correction mechanism can be integrated in overlapping based FL protocols by using stale averaged gradients for homogeneous FL client devices at the huge cost of their memory storage [9]. Heterogeneity among clients’ mobile devices makes the memory cost issue more severe and almost diminishes the benefits of overlapping computing and communication. For example, following typical overlapping based FL protocols, fast devices (computing + communications) have to wait for the slowest device (i.e., straggler) to finish an FL training round, and those fast devices, in particular fast computing devices, will undergo many rounds of local training in advance before receiving the new/next round global model for update correction. That will accumulate rounds of gradients to correct, causing significant model staleness, and impose a huge burden on, or even overflow fast mobile devices’ memory. The advanced local training of those fast devices can also drive local models towards their respective local optima, potentially resulting in divergence from the objective of the global model (a.k.a model drift [10]) and hindering FL convergence.
@@
Figure 1: Overlapping deficiencies due to heterogeneity.
To alleviate the straggler issue in heterogeneous environments, researchers have explored to select a subset of devices to participate in FL training over rounds, i.e., FL participation selection (PS) [11, 12, 13, 14]. For example, Oort [15], a pioneering FL PS design, jointly considers the statistical utility (i.e., FL learning convergence/accuracy) and global system utility (i.e., training latency per round). Oort selects participating mobile devices that have good contributions to FL learning convergence and are capable of completing local training before a setup latency threshold. However, existing FL PS methods still follow traditional computing and communication interleaving FL training protocol and don’t consider the training speedup brought by continuous computing or overlapping.
@@
  • We develop a FedEx prototype and evaluate its performance with extensive experiments. The experimental results demonstrate that FedEx can remarkably reduce the latency for FL training over heterogeneous mobile devices at limited costs. Our code has been open-sourced at: https://github.com/BESTTOOLBOX/FedEx.
TABLE I: Comp. vs Comm. (Squeezenet@CIFAR10).
Mobile Device Configuration [^1] The transmission data rates displayed in Table I represent the averaged achievable data rates for uploading transmissions. Figure 2 employs the same communication settings. $T_{n}^{cp}$ $T_{n}^{cm}$ $T_{n}^{cm}/T_{n}^{cp}$
Xavier w/ Wi-Fi (6.9Mbps) 1.13s 5.54s 4.90
TX2 w/ Wi-Fi (6.0Mbps) 1.35s 6.40s 4.74
Xiaomi 12S w/ LTE (5.0Mbps) 0.84s 7.66s 9.12
II Background and Motivations
II-A FL on Mobile Devices
@@
$\min_{w\in\mathbb{R}^{d}}F\left(w\right):=\sum_{n=1}^{N}\pi_{n}F_{n}\left(w\right)\;,$ (1)
@@
II-B Deficiencies of Overlapping Based FL in Heterogeneous Environments
@@
Figure 2: Performance comparison: homo. vs heter. (CNN@MNIST, non-i.i.d. data).
Specifically, following DGA procedure [^2] Please refer to “Algorithm 1 Delayed Gradient Averaging (DGA)” in [9]., the device transmits its local updates when it completes every fixed number of local training iterations, say $K$ iterations. Nevertheless, it may lead to “collision of local update communications” issues, if certain device’s transmission latency is much larger than its computing time consumption as illustrated in Fig. 1(a). Even if the mobile device can be equipped with multiple antennas allowing for simultaneous transmissions of local updates over different rounds, it will transmit the stale local model gradients, i.e., locally trained based on the ($r-1$)-th round global model, for the ($r+1$)-th round updates before it receives the $r$-th round global model for staleness correction. We conducted empirical studies with popular mobile devices and accessing technologies, and the results in Table I show that the fast computing and slow transmission situations largely exist in practice, which can trigger the collision of communications for local model updates.
Assuming the overlapping training can proceed without “collision of local model updates” in heterogeneous environments, the memory cost issue for model update correction will be aggravated due to computing and wireless heterogeneity among mobile devices and overlapping model staleness may incur model drift issue due to data heterogeneity, which can degrade or even fail FL convergence. Overlapping based FL requires each device to locally store multiple copies of its old local model updates for update correction, and the memory consumption of the mobile device increases linearly with the staleness level. Different from homogeneous environments, computing and wireless heterogeneity enlarges the time consumption gap between fast devices and the slowest device (i.e., straggler) to complete their local computing and model updates. The enlarged gap directly increases the staleness level of fast devices. When the overlap staleness $S_{n}$ is large, the heavy gradient storage requirement will impose a huge burden on, or even overflow fast mobile devices’ limited memory. Moreover, due to data heterogeneity, the local trained models point to the direction towards the local optima, which may not be consistent with the objective of the aggregated global model, called model drift [10]. Thus, a high overlapping staleness level, i.e., performing too much continuous local computing, is detrimental to the learning accuracy and may significantly slow down or fail FL convergence under data heterogeneous scenarios. Fig. 2 shows our empirical studies to compare the testing accuracy of DGA and FedAvg in homogeneous and heterogeneous environments, respectively, which supports our model drift and staleness analysis above.
@@
II-C Limitations of SOTA PS Methods
To mitigate the straggler issue in heterogeneous environments, various FL PS designs have been proposed to select a subset of devices to participate in FL training over rounds [11, 12, 13, 14]. As an advanced FL PS design, Oort [15] novelly considers both the statistical utility (i.e., FL learning convergence/accuracy) and global system utility (i.e., training latency per round). The utility function of device $n$ in Oort can be expressed as follows.
$\small\begin{split}Util(n)=|B_{n}|\sqrt{\frac{1}{|B_{n}|}\sum_{\begin{subarray}{c}k\in{B}_{n}\end{subarray}}Loss(k)^{2}}\times(\frac{T}{t(n)})^{\mathbb{I}(T<t(n))\times\alpha}\;,\end{split}\vspace{-2mm}$ (2)
where $B_{n}$ represents the local training samples of mobile device $n$, $Loss(k)$ corresponds to the training loss of the $k$-th sample, $t(n)$ signifies the system latency of mobile device $n$, $T$ is the developer-preferred duration for each round, and $\alpha$ serves as the penalty factor. Additionally, $\mathbb{I}(x)$ denotes an indicator function, where $\mathbb{I}(x)=1$ when $x$ is true, and 00 otherwise. Under Eqn. (2), top $K$ devices with the highest utility will be selected for FL training in synchronous manner. Instead of picking devices with top-$K$ utility deterministically over rounds, Oort employs a “temporal uncertainty” mechanism [15] to gradually increase the utility of a device if it has been overlooked for a long time. Thus, low-utility devices still have chances to be selected to avoid biased FL training.
@@ II-D Challenges in Overlapping + PS
Secondly, PS utility function doesn’t take the latency reduction brought by overlapping into account. Simply, the overlapping and continuous training changes the way of traditional computing and communication interleaved FL training, which consequently affects the definition of computing latency in a training round and corresponding per round training latency for a candidate mobile device. While the global system utility [15] in Oort’s utility function is based on conventional per round FL training latency, directly using Oort to select the participating mobile devices for overlapping based FL is not appropriate, since it is not aware of overlapping benefits to accelerate FL training convergence.
@@
III FedEx Design
III-A FedEx Overview
@@
Figure 3: The sketch of FedEx procedure and testbed.
III-B Model Updates with Staleness Ceiling
To preclude the collision of communications, alleviate model drift, constrain the memory cost for continuous computing (Sec. II-B) and synchronize devices’ local model updates in every round (Sec. II-D), we propose a novel local model updates protocol with staleness ceiling in FedEx. As discussed in Sec. II, we found that in heterogeneous environments, existing overlapping principle (e.g. DGA) of uploading local gradients after every $K$ computing iterations is the root for the collision of communications and asynchronous model updates, and its endless/non-constrained continuous local computing aggravates model staleness, which results in model drift as well as huge memory consumption for updates correction. So, to address those issues and fit FL training over heterogeneous mobile devices, FedEx redefines the rule of overlapping local model updates and casts a ceiling on staleness/the iteration number of continuous local computing.
@@ III-B Model Updates with Staleness Ceiling
III-C Overlapping Aware PS Utility Function
@@
To address straggler issues for latency reduction per round (Sec. II-B) and make PS aware of overlapping benefits (Sec. II-C and Sec. II-D), we develop an overlapping aware utility function for FL PS in heterogeneous environments. Inspired by Oort’s utility function [15] in Eqn. (2), the proposed PS utility function aims to jointly improve statistical utility and delay efficiency of FL training, while considering the latency reduction brought by overlapping computing and communication. Thus, we revise global system utility in Eqn. (2) into overlapping latency utility in FedEx, and present the overlapping aware PS utility function as follows.
@@
$\vspace{-3mm}\small\begin{split}U(n,r)\!=\!\underbrace{|B_{n}^{r}|\sqrt{\frac{1}{|B_{n}^{r}|}\sum_{\begin{subarray}{c}k\in{B}_{n}^{r}\end{subarray}}\!Loss(k)^{2}}}_{\textbf{Statistical utility}}\!\times\!\underbrace{\left(\frac{1}{(K\!-\!S_{n}^{r-1})\!\cdot\!t_{n}^{r}\!+\!\frac{D_{n}}{R_{n}^{r}}}\right)^{\alpha}}_{\textbf{Overlapping latency utility}},\end{split}\vspace{-3mm}$ (3)
@@
where $|B_{n}^{r}|$ is the number of training data samples of mobile device $n$ in the $r$-th round, $t_{n}^{r}$ is the averaged time consumption for one iteration of local computing at mobile device $n$ [16], $D_{n}$ is the model size of client $n$ and $R_{n}^{r}$ is the transmission rate of mobile device $n$ in the $r$-th round. $\frac{D_{n}}{R_{n}^{r}}$ represents the mobile device $n$’s estimated transmission time. Here, $\alpha$ is the scaling coefficient to trade off the statistical utility and overlapping latency utility.
@@
The utility function above consists of two parts: statistical utility and overlapping latency utility. The statistical utility accounts for the estimated learning performance contributions to FL convergence, which is the same as that in Oort [15]. The overlapping latency utility has two components: classical computing latency and the model update transmission latency. For classical computing latency in the $r$-th round, mobile device $n$ needs to complete $K$ local training iterations. Regarding the latency reduction by overlapping, device $n$ has already conducted continuous computing for $S_{n}^{r-1}$ iterations in the ($r-1$)-th round. Thus, in the $r$-th round, device $n$ only needs to perform ($K-S_{n}^{r-1}$) iterations for classical computing phase before transmitting its local gradients. As illustrated in [16], given a learning task and the computing capability of mobile device $n$, $t_{n}^{r}$ is relatively fixed in practice and can be estimated with good precision by averaging the latency for the previous iterations. Therefore, device $n$’s classical computing latency in the $r$-th round can be estimated as $(K-S_{n}^{r-1})\cdot t_{n}^{r}$. For transmission latency, given the fact that mobile devices’ wireless communication conditions don’t significantly change within a short period of time, we assume $R_{n}^{r}$ is fixed in the $r$-th round and can be approximated by using $n$’s transmission rate at the beginning of this round.
@@
The proposed overlapping latency utility is aware of overlapping benefits by considering latency reduction $(K-S_{n}^{r-1})\cdot t_{n}^{r}$. Besides, as discussed in Sec. III-B, the per-round latency in FedEx is determined by the straggler’s training latency, i.e., the slowest device’s classical on-device training time plus its communication time for transmitting local updates. The overlapping latency utility helps to address straggler issues since it will select fast devices with small training latency in the round, i.e., $[(K-S_{n}^{r-1})\cdot t_{n}^{r}+\frac{D_{n}}{R_{n}^{r}}]$. Together with statistical utility, the proposed overlapping aware utility function [^3] Regarding the privacy preservation of statistical utility and overlap latency utility, we leverage the same privacy analysis as presented in Oort [15]. will select the mobile devices with good learning contributions and high overlapping latency utility to participate, which can accelerate the FL training over heterogeneous mobile devices [^4] For device $n$ not selected to participate in the $r$-th round, it refrains from any computing and transmission activities. Instead, it inherits the values of $K$ and $S_{n}^{r-1}$ from the $(r-1)$-th round, updates them into $K$ and $S_{n}^{r}$, and wait for the next time to participate. For devices overlooked for a long time, we follow the “temporal uncertainty” mechanism to gradually increase their utilities, in order to avoid biased training in heterogeneous environments. Due to page limits, the duplicated discussions have been omitted here..
III-D FedEx Overlapping Trigger